40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 14 07:51:20 2011, steve, Update, SUS, PRM damping restored 
    Reply  Thu Sep 15 07:28:08 2011, steve, Update, SUS, PRM damping restored 
Message ID: 5415     Entry time: Thu Sep 15 07:28:08 2011     In reply to: 5403
Author: steve 
Type: Update 
Category: SUS 
Subject: PRM damping restored 

Quote:

The PRM damping was restored at side sensor var 1050

 The PRM sus damping restored.

ELOG V3.1.3-